# A concise tutorial review of Reverse Osmosis and Electrodialysis

### v5

@@ Abstract
Reverse osmosis (RO) and electrodialysis (ED) are the two most important membrane technologies for water desalination and treatment. Their modes of operation and transport mechanisms are very different, but on a closer look also have many similarities. In this concise version of our tutorial review, we describe state-of-the-art theory for both processes, focusing on simple examples that are helpful for the non-specialist and useful for classroom teaching. Both processes are described by the solution (SF) theory which combines ion and water transport across membranes with chemical and mechanical equilibrium at membrane/solution interfaces. We present a derivation of SF theory based on force balances on water and ions and show how the various terms, convection, diffusion, and electromigration, are derived, and how solute partitioning is implemented. Finally, we demonstrate how SF theory accurately describes the osmosis experiment where water and ions are transported in opposite directions across a membrane.
@@ 1 Introduction
In this tutorial we focus on desalination of water, i.e., how to use membranes to obtain freshwater from brackish water or seawater by RO and ED. In section 2 we describe thermodynamic equations and metrics that apply to both methods. In sections 3–5 we present the solution-friction (SF) theory, which describes ion and water flow through membranes in combination with chemical and mechanical equilibrium at solution/membrane interfaces. In section 3 we use this theory for neutral solutes and present analytical solutions for solute rejection, also including the effect of concentration polarization (CP). For ED we discuss in section 4 a co-current plug flow model also based on SF theory but in the absence of water flow across the membranes. In section 5 we provide a fundamental derivation of SF theory both for ion transport and for water transport, and we subsequently apply SF theory to describe data of an osmosis experiment where we have counter-directional salt and water flow. We also discuss the origin of osmosis, i.e., the reason why water flows from a solution at low salinity to a solution of high salinity. Section 6 is conclusions and outlook.
We illustrate in Fig. 2 that RO and ED have many aspects in common, which they also have in common with the phenomenon of osmosis. In all these processes, the same three features play a key role, as we discuss next. First, in all cases there is a difference in concentration of solutes, or salt, between the two sides of the membrane. Furthermore, there is a certain electrical current, and finally there is a pressure difference across the membrane. In certain cases, the electrical current will be zero, as in RO, and likewise the hydrostatic pressure difference will be (close to) zero in some cases, for instance in the osmosis experiment when the two sides are open to atmosphere and the liquid levels about equal. Nevertheless, in the transport theory, current and pressure being zero is part of the model structure. And thus, when building a theoretical model, it does not matter very much whether current is zero or non-zero, and in all cases a relation for current is part of the model structure. And we have the same situation for the pressure difference. Because of these similarities in the required theoretical description for RO, ED, and osmosis, a combined discussion of these processes is possible.
@@ 2.1 Overall mass balances for water desalination
where $\phi_{\text{s},j}$ is a volumetric flow rate, for instance in m3/hr. Secondly there is a mass balance per solute type
@@ 2.2 Setpoints and metrics for water desalination
Every water desalination process is defined by certain operational characteristics, i.e., objectives, or setpoints. These are the amount of water to be treated, the water recovery, and the extent of desalination (or salt removal). Different setpoints can be defined but in the end they describe how much freshwater is obtained from how much feedwater, and what is the salt concentration of the product water. These are considered as setpoints and are ‘hard’ objectives that must be attained.
After the setpoints of a process have been defined, metrics (or, equivalently, ‘performance indicators’) come into play. They describe optimized operation. Many metrics are possible. We can for instance define a measure of energy use (energy efficiency), the degree of removal of specific compounds (selectivity), or a measure of how long membranes remain functional.
@@ 2.2.1 Setpoints for water desalination
Setpoints for a water desalination process are operational parameters that are ‘hard’, that must be attained. It is not the case that they are just ‘nice to have’, i.e., to be optimized in some direction. Instead, they define a certain desalination operation. They are also essential for a calculation of the theoretical minimum energy use (see section 2.3). In practice, the aim is not that these setpoints are reached in an individual desalination unit (module, stack), but they will be the result of operation of a total desalination plant. So a first module desalinates to a certain percentage, and only in a second module the final salt concentration is reached. In that case, setpoints are only reached at the overall level of a full water desalination plant.
@@
Other terminology for water recovery is recovery ratio, and other symbols that are used are r, RR, or $\alpha$. A related definition is ‘split ratio’ for the ratio $\phi\mathrm{{}_{v,p}}{\big{/}}\phi\mathrm{{}_{v,c}}$, which is equal to $\textsc{wr}/\left(1-\textsc{wr}\right)$.
@@ 2.2.2 Metrics for water desalination
A first metric tracks the specific removal of certain key compounds, and the more the removal of these, the better. So a metric could be a removal degree of one or several key ions, perhaps expressed as single number by summing over several key ions including a weighing factor describing the importance of the removal of each.
Another example of a metric is the energy efficiency, $\eta$, which is the ratio of the theoretical minimum energy to achieve a certain desalination (defined by the setpoints) over the actual energy investment. (This metric is also called TEE, for ‘thermodynamic energy efficiency’.) There are many possible definitions for $\eta$, depending on which energy contributions are included in the actual energy input. For instance, for ED we can limit analysis to the electrical energy input required for current to flow across the stack, but we can also include the (much smaller) costs of pumping the water through the channels. In RO we can decide to subtract the recovery of ‘pressure energy.’ We can take into account efficiencies of power sources and pumps, and of other electric equipment, for instance to convert (AC) electricity into pumping power for RO, or into DC stack current for ED [14]. So, a report detailing values of $\eta$ must carefully lay out which energy factors are considered.
Often, instead of an efficiency $\eta$, the inverse factor is reported, which then is the factor by which the actual desalination energy is ‘above the minimum.’ Generally, values as low as two, for this inverse factor, are considered to be very low, i.e., we now have an actual energy use that is only a factor two above the thermodynamic minimum. In many cases, especially for water of low salinity, this factor is much higher, for instance 5 or 10, which implies the actual energy is equal to that number of times the thermodynamic minimum.
@@ 2.3.2 Extensions for non-ideal solutions
The first ‘non-ideal’ effect is that of electrostatic Coulombic interactions between ions, which for a 1:1 salt can be described by the extended Bjerrum equation [15, 16]
${f_{\text{elec}}}=-{2RTc_{\infty}}\cdot\left(\alpha_{1}\cdot\overline{c}^{1/3}+\alpha^{2}\cdot\overline{c}^{2/3}-\alpha_{3}\cdot\overline{c}\right)$ (7)
@@
As a numerical example, for a 1:1 salt at $c_{\infty}\!=\!525$ mM, with ions of a hydrated size of $\sigma\!=\!0.5$ nm (size is twice the radius), the ion volume fraction is $\phi\!=\!0.0414$. With the electrostatic and volumetric contributions implemented, then for the previous numerical example of a 50% water recovery, the minimum energy consumption increases from $E\mathrm{{}_{min}}$=1.00 kWh/m3 to $E\mathrm{{}_{min}}$=1.23 kWh/m3. Thus, with the corrections included, the minimum energy increases, and thus for a given achieved energy use in practice, the efficiency is better than in case the ideal value for $E\mathrm{{}_{min}}$ was assumed. Eq. (8) can be extended to describe a salt solution where the ions have different sizes. Further extensions relate to the energies of ion-ion association and ion (de-)protonation, with pK-values of these reactions dependent on salt concentration, but this is more complicated.
@@ 2.3.3 Simplifications of the thermodynamic expressions
When we combine the various balances with the expressions for $S_{i}$, we obtain for $E\mathrm{{}_{min}}$, the result that[18]
@@
which shows that the minimum required energy goes up when we aim for a higher water recovery. The factor ‘2’ on the left side here, is because we consider a $z$:$z$ salt. For a single neutral solute, not a $z$:$z$ salt, then factors ‘2’ on the left sides of Eqs. (9) and (10) must be omitted.
Note that Eq. (10) is only correct when $c\mathrm{{}_{p}}=0$. For non-zero values of $c\mathrm{{}_{p}}$, this energy, for very small wr, is given by
@@ 2.4 Practical energy use in RO and ED
The energy required to pressurize a stream is $\Delta P^{\text{h}}$ times flowrate $\phi_{\text{v}}$, which is a number with unit J/s=W. This energy we need to pump feedwater into a module. And to pump water exiting one module into the next one, if it operates at a higher pressure. At the end of a series of modules, it is possible to install an energy recovery device (ERD). If the water on the retentate side (the concentrate; a smaller volume than the flow of feedwater) passes an ERD, exiting at a lower pressure, energy recovery is possible, given by $\eta_{\mathrm{{}_{ERD}}}\cdot\Delta P^{\text{h}}\cdot\phi_{\text{v,c}}$, with $\eta_{\mathrm{{}_{ERD}}}$ the energy efficiency of the ERD.
Another practical limitation, even for ideal operation, is that we generally have a limited number of modules. In RO with a single module, for a membrane that perfectly retains all solutes, the minimum pressure is given by the retentate (concentrate) concentration, which in case of perfect solute rejection relates to $c\mathrm{{}_{f}}$ by $c\mathrm{{}_{c}}=c\mathrm{{}_{f}}/\left(1-\text{{wr}}\right)$. So the minimum energy per unit permeate volume, practically, cannot be less than $E\mathrm{{}_{min,prac}}=RT\,c\mathrm{{}_{f}}\,/\left(1-\text{{wr}}\right)$ [18], and with $E\mathrm{{}_{min}}$ given by Eq. (10), this implies a practical single stage RO efficiency of
@@ 3 Theory of reverse osmosis with neutral solutes
3.1 Flowrates and rejection in a 1D model of RO for neutral solutes
Reverse osmosis (RO) is a process to remove solutes from water by pushing water through a membrane which largely blocks passage for solutes. This is the general mechanism of all pressure-driven membrane processes, but for water desalination the size of the pores (free volume) in the membrane must be very small, typically well below 1 nm in the selective toplayer of an RO membrane. In classical literature, RO is called hyperfiltration, but nowadays this term has been replaced by RO. Nanofiltration (NF) is similar to RO but has slightly larger pores, and thus the applied hydrostatic pressure is lower, with the rejection of monovalent ions significantly reduced. NF is therefore considered to be suitable for the selective separation of divalent ions and other larger molecules from monovalent salts.
@@ 3.1 Flowrates and rejection in a 1D model of RO for neutral solutes
$J_{i}=K_{\text{f},i}\,c_{\text{m},i}v\mathrm{{}_{w}}-K_{\text{f},i}D_{\text{m},i}\frac{\partial c_{\text{m},i}}{\partial x}$ (13)
where the transmembrane solute flux $J_{i}$ (in mol/m2/s) and water flux $v\mathrm{{}_{w}}$ (in m/s) are defined per unit geometric membrane area, thus they are superficial velocities, i.e., they are not pore-based, interstitial, velocities. Concentrations $c_{\text{m},i}$ in the membrane are defined per unit volume of the water-filled pores in the membrane. Eq. (13) includes solute transport due to convection, which is the transport of solutes because they ‘flow with’, i.e., are convected with the water that flows through the membrane (first term), and transport because of a solute concentration gradient (second term). Here, x is a coordinate axis directed across the membrane. The diffusional term depends on a membrane-based diffusion coefficient, $D_{\text{m},i}$, which includes the effects of porosity and tortuosity. The friction factor $K_{\text{f},i}$ depends on the extent of solute-membrane friction, and thus relates to membrane pore size and the (hydrated) size of the solutes. In the absence of solute-membrane friction, the solutes move freely with the water in the pores, while they are still subject to diffusional forces (and electrical forces when charged), and in that case $K_{\text{f},i}\!=\!1$. However, in an RO process, it is much more likely that $K_{\text{f},i}\ll 1$.
@@
$\Phi_{i}=S_{i}=\frac{c_{\text{m},i}}{c_{\infty,i}}$ (14)
where index ‘m’ refers to a position just in the membrane and $\infty$’ to a position just outside the membrane. When the solutes are charged, i.e., they are ions, an additional Donnan effect arises, which in this paper we treat separately, i.e., do not include in $S_{i}$ or $\Phi_{i}$, see section 5.2. The partitioning coefficient, $\Phi_{i}$, is the same as the solubility of a species in a material, $S_{i}$, and this term is for instance due the chemical affinity of a solute with a certain phase (medium) relative to its affinity with another phase, but also other effects of solute-membrane interaction can be absorbed in $\Phi_{i}$ [22, 23, 16]. In RO with neutral solutes, we always have at least $\Phi_{i}\!<\!1$ or $K_{\text{f},i}\!<\!1$, but likely both factors are $<\!1$.
A further element in RO modeling is the ‘permeate equation’, an equation essential to describe RO in a one-dimensional (1D) geometry, valid when there is no flow along the membrane on the permeate side, which is Eq. (8) in ref. [5], and Eq. (24) in ref. [24], given by
$c_{\text{p},i}=\frac{J_{i}}{v\mathrm{{}_{w}}}$ (15)
which looks very simple, but is actually a quite intriguing result. It states that the concentration of solutes on the permeate side is a direct function of the ratio of solute flux through the membrane over the water flux. This is a correct result in steady-state, thus with conditions on the upstream side unvarying in time. It is also valid when on the upstream side conditions change in time and the permeate is regularly removed; as a consequence water that permeated the membrane at earlier moments does not mix up with fresh permeate. In a complete (2D) module, Eq. (15) does not apply at each position along the membrane, but it can be used to describe the exit permeate flow, when we use average values for $J_{i}$ and $v\mathrm{{}_{w}}$ (each separately averaged over the complete module).
@@
${J_{i}}=\left(1-\sigma_{i}\right)\,v\mathrm{{}_{w}}\,\frac{c_{\text{f},i}\cdot e^{\text{Pe}_{i}}-c_{\text{p},i}}{e^{\text{Pe}_{i}}-1}$ (16)
where $\text{Pe}_{i}$ is the membrane Péclet number, $\text{Pe}_{i}=v\mathrm{{}_{w}}/k_{\text{m},i}$, where $k_{\text{m},i}$ is the membrane mass transfer coefficient, given by $k_{\text{m},i}=D_{\text{m},i}{\big{/}}L\mathrm{{}_{m}}$, with Lm membrane thickness. In Eq. (16), $\sigma_{i}$ is the sieving coefficient which relates to $K_{\text{f},i}$ and $\Phi_{i}$ according to
@@
Eq. (16) is the general result for rejection in an RO membrane with neutral solutes [24, 28, 26, 27, 25, 5]. Because of the dependence of solubility, i.e., partitioning, described by the factor $\Phi_{i}$, in combination with a derivation based on ion-water and ion-membrane friction, captured by the factor $K_{\text{f},i}$, we call this the solution-friction model, or SF-model. Calculation results based on this equation are presented in Fig. 3. Note that when concentration polarization (CP) is also considered, Eq. (16) is still valid, but we must replace $c_{\text{f},i}$ by the concentration right at the membrane interface, often written as $c_{\text{int},i}$.
@@
which is equivalent to Eq. (4) in ref. [20], Eq. (3) in ref. [27], Eq. (38) in ref. [29], and Eq. (129) in ref. [30], when we implement that $\omega=\left(1-\sigma_{i}\right)\,k_{\text{m},i}$, $P=P\mathrm{{}_{s}}=K_{\text{f}}\Phi_{i}k_{\text{m},i}$, and thus $k_{\text{m},i}=P\mathrm{{}_{s}}/\left(1-\sigma_{i}\right)$. A similar formulation is Eq. (9) in ref. [31]. Note that while Eq. (16) is generally valid, at each position in an RO module, Eq. (18) is only valid in the limit of a low water recovery. All equations in this section are derived for neutral solutes.
Based on Eq. (18), rejection in the limit of low transmembrane water flux $v\mathrm{{}_{w}}$, is given by
$R_{i}=\frac{\sigma_{i}}{1-\sigma_{i}}\;\text{Pe}_{i}=\frac{\sigma_{i}}{\omega}\;v\mathrm{{}_{w}}$ (19)
which shows that for $v\mathrm{{}_{w}}\rightarrow 0$, the solute rejection goes to zero, and it also shows hat to have a non-zero rejection we need $\sigma_{i}>0$, and thus we either need a non-unity $K_{\mathrm{f},i}$-factor or a non-unity solubility, $\Phi_{i}$.
In the other limit, of a high permeate water flux, $v\mathrm{{}_{w}}$, Eq. (16) simplifies to
$J_{i}=K_{\text{f},i}\Phi_{i}\,c_{\text{f},i}\,v\mathrm{{}_{w}}=\left(1-\sigma_{i}\right)\,c_{\text{f},i}\,v\mathrm{{}_{w}}$ (20)

### v6

@@ Abstract
Reverse osmosis (RO) and electrodialysis (ED) are the two most important membrane technologies for water desalination and treatment. Their modes of operation and transport mechanisms are very different, but on a closer look also have many similarities. In this concise version of our tutorial review, we describe state-of-the-art theory for both processes, focusing on simple examples that are helpful for the non-specialist and useful for classroom teaching. Both processes are described by solution-friction (SF) theory which combines ion and water transport across membranes with chemical and mechanical equilibrium at membrane/solution interfaces. We present a derivation of SF theory based on force balances on water and ions and show how the various terms, convection, diffusion, and electromigration, are derived, and how solute partitioning is implemented. Finally, we demonstrate how SF theory accurately describes the osmosis experiment where water and ions are transported in opposite directions across a membrane.
@@ 1 Introduction
In this tutorial we focus on desalination of water, i.e., how to use membranes to obtain freshwater from brackish water or seawater by RO and ED. In section 2 we describe thermodynamic equations and metrics that apply to both methods. In sections 3–5 we present the solution-friction (SF) theory, which describes ion and water flow through membranes in combination with the assumption of chemical and mechanical equilibrium at solution/membrane interfaces. In section 3 we use this theory for neutral solutes and present analytical solutions for solute rejection, also including the effect of concentration polarization (CP). For ED we discuss in section 4 a co-current plug flow model also based on SF theory but in the absence of water flow across the membranes. In section 5 we provide a fundamental derivation of SF theory both for ion transport and for water transport, and we subsequently apply SF theory to describe data of an osmosis experiment where we have counter-directional salt and water flow. We also discuss the origin of osmosis, i.e., the reason why water flows from a solution at low salinity to a solution of high salinity. Section 6 is conclusions and outlook.
We illustrate in Fig. 2 that RO and ED have many aspects in common, which they also have in common with the phenomenon of osmosis. In all these processes, the same three features play a key role, as we discuss next. First, in all cases there is a difference in concentration of solutes, or salt, between the two sides of the membrane. Furthermore, there is a certain electrical current, and finally there is a pressure difference across the membrane. In certain cases, the electrical current will be zero, as in RO, and likewise the hydrostatic pressure difference will be (close to) zero in some cases, for instance in the osmosis experiment when the two sides are open to atmosphere and the liquid levels about equal. Nevertheless, in the transport theory, current and pressure being zero is part of the model structure. And thus, when building a theoretical model, it does not matter very much whether current is zero or non-zero, and in all cases a relation for current is part of the model structure. And we have the same situation for the pressure difference (unless in ED we assume water flow across the membrane is zero). Because of these similarities in the required theoretical description for RO, ED, and osmosis, a combined discussion of these processes is possible.
@@ 2.1 Overall mass balances for water desalination
where $\phi_{\text{v},j}$ is a volumetric flow rate, for instance in m3/hr. Secondly there is a mass balance per solute type
@@ 2.2 Setpoints and metrics for water desalination
Every water desalination process is defined by certain operational characteristics, i.e., objectives, or setpoints. These are the amount of water to be treated, the water recovery, and the extent of desalination (or salt removal). Different setpoints can be defined but in the end they describe how much freshwater is obtained from how much feedwater, and what is the salt concentration of the product water. These are considered as setpoints and they must be attained.
After the setpoints of a process have been defined, metrics, or, equivalently, performance indicators, come into play. They describe optimized operation. Many metrics are possible. We can for instance define a measure of energy use (energy efficiency), the degree of removal of specific compounds (selectivity), or a measure of how long membranes remain functional.
@@ 2.2.1 Setpoints for water desalination
Setpoints for a water desalination process are operational parameters that must be attained, because they define a certain desalination operation. They are also essential for a calculation of the theoretical minimum energy use (see section 2.3). In practice, the aim is not that these setpoints are reached in an individual desalination unit (module, stack), but they will be the result of operation of a total desalination plant. So a first module desalinates to a certain percentage, and only in a second module the final salt concentration is reached. In that case, setpoints are only reached at the overall level of a full water desalination plant.
@@
Other terminology for water recovery is recovery ratio, and other symbols that are used are r, RR, or $\alpha$. A related definition is split ratio for the ratio $\phi\mathrm{{}_{v,p}}{\big{/}}\phi\mathrm{{}_{v,c}}$, which is equal to $\textsc{wr}/\left(1-\textsc{wr}\right)$.
@@ 2.2.2 Metrics for water desalination
A first metric tracks the specific removal of certain key compounds, and the higher the removal of these, the better. So a metric could be a removal degree of one or several key ions, perhaps expressed as a single number by summing over several key ions including a weighing factor describing the importance of the removal of each.
Another example of a metric is the energy efficiency, $\eta$, which is the ratio of the theoretical minimum energy to achieve a certain desalination (defined by the setpoints) over the actual energy investment. This metric is also called TEE, for thermodynamic energy efficiency. There are many possible definitions for $\eta$, depending on which energy contributions are included in the actual energy input. For instance, for ED we can limit analysis to the electrical energy input required for current to flow across the stack, but we can also include the (much smaller) costs of pumping the water through the channels. In RO we can decide to subtract the recovery of pressure energy. We can take into account efficiencies of power sources and pumps, and of other electric equipment, for instance to convert (AC) electricity into pumping power for RO, or into direct current for ED [15]. So, a report detailing values of $\eta$ must carefully lay out which energy factors are considered.
Often, instead of an efficiency $\eta$, the inverse factor is reported, which then is the factor by which the actual desalination energy is above the minimum. Generally, values as low as two, for this inverse factor, are considered to be very low, i.e., we now have an actual energy use that is only a factor two above the thermodynamic minimum. In many cases, especially for water of low salinity, this factor is much higher, for instance 5 or 10, which implies the actual energy is equal to that number of times the thermodynamic minimum.
@@ 2.3.2 Extensions for non-ideal solutions
The first non-ideal effect is that of Coulombic interactions between ions, which for a 1:1 salt can be described by the extended Bjerrum equation [16, 17]
${f_{\text{cou}}}=-{2RTc_{\infty}}\,\left(\tfrac{3}{4}b{c_{\infty}}^{1/3}+\tfrac{3}{20}b^{2}{c_{\infty}}^{2/3}-3b^{3}q{c_{\infty}}\right)$ (7)
@@
As a numerical example, for a 1:1 salt at $c_{\infty}\!=\!525$ mM, with ions of a hydrated size of $\sigma\!=\!0.5$ nm (size is twice the radius), the ion volume fraction is $\phi\!=\!0.0414$. With the Coulombic and volumetric contributions implemented, then for the previous numerical example of a 50% water recovery, the minimum energy consumption increases from $E\mathrm{{}_{min}}$=1.00 kWh/m3 to $E\mathrm{{}_{min}}$=1.23 kWh/m3. Thus, with the corrections included, the minimum energy increases, and thus for a given achieved energy use in practice, the efficiency is better than in case the ideal value for $E\mathrm{{}_{min}}$ was assumed. Eq. (8) can be extended to describe a salt solution where ions have different sizes. Further extensions relate to the energies of ion-ion association and ion (de-)protonation, with pK-values of these reactions dependent on salt concentration, but this is more complicated.
@@ 2.3.3 Simplifications of the thermodynamic expressions
When we combine the various balances with the expressions for $S_{i}$, we obtain for $E\mathrm{{}_{min}}$ the result that[19]
@@
which shows that the minimum required energy goes up when we aim for a higher water recovery. Note that Eq. (10) is only correct when $c\mathrm{{}_{p}}=0$. For non-zero values of $c\mathrm{{}_{p}}$, this energy, for a very small wr, is given by
@@
The factor 2 on the left of Eqs. (9)–(11) is because we consider here a $z$:$z$ salt. For a single neutral solute, this factor is omitted.
@@ 2.4 Practical energy use in RO and ED
The energy required to pressurize a stream is $\Delta P^{\text{h}}$ times flowrate $\phi_{\text{v}}$, which is a number with unit J/s=W. This energy we need to pump feedwater into a module, and to pump water exiting one module into the next one, if it operates at a higher pressure. At the end of a series of modules, it is possible to install an energy recovery device (ERD). If the water on the retentate side (the concentrate; a smaller volume than the flow of feedwater) passes an ERD, exiting at a lower pressure, energy recovery is possible, given by $\eta_{\mathrm{{}_{ERD}}}\cdot\Delta P^{\text{h}}\cdot\phi_{\text{v,c}}$, with $\eta_{\mathrm{{}_{ERD}}}$ the energy efficiency of the ERD.
Another practical limitation, even for ideal operation, is that we generally have a limited number of modules. In RO with a single module, for a membrane that perfectly retains all solutes, the minimum pressure is given by the retentate (concentrate) concentration, which in case of perfect solute rejection relates to $c\mathrm{{}_{f}}$ by $c\mathrm{{}_{c}}=c\mathrm{{}_{f}}/\left(1-\text{{wr}}\right)$. So the minimum energy per unit permeate volume, practically, cannot be less than $E\mathrm{{}_{min,prac}}=2\,c\mathrm{{}_{f}}\,RT/\left(1-\text{{wr}}\right)$ [19], and with $E\mathrm{{}_{min}}$ given by Eq. (10), this implies a practical single stage RO efficiency of
@@ 3 Theory of reverse osmosis with neutral solutes
3.1 Flowrates and rejection
Reverse osmosis (RO) is a process to remove solutes from water by pushing water through a membrane which largely blocks passage for solutes. This is the general mechanism of all pressure-driven membrane processes, but for water desalination the size of the pores (free volume) in the membrane must be very small, typically below 1 nm in the selective toplayer of an RO membrane. In classical literature, RO is called hyperfiltration, but nowadays this term has been replaced by RO. Nanofiltration (NF) is similar to RO but has slightly larger pores, and thus the applied hydrostatic pressure is lower, with the rejection of monovalent ions significantly reduced. NF is therefore considered to be suitable for the selective separation of divalent ions and other larger molecules from monovalent salts.
@@ 3.1 Flowrates and rejection
$J_{i}=K_{\text{f},i}\,c_{\text{m},i}J\mathrm{{}_{w}}-K_{\text{f},i}D_{\text{m},i}\frac{\partial c_{\text{m},i}}{\partial x}$ (13)
where the transmembrane solute flux $J_{i}$ (in mol/m2/s) and water flux $J\mathrm{{}_{w}}$ (in m/s) are defined per unit geometric membrane area, thus they are superficial velocities, i.e., they are not pore-based, interstitial, velocities. Concentrations $c_{\text{m},i}$ in the membrane are defined per unit volume of the water-filled pores in the membrane. Eq. (13) includes solute transport due to convection, which is the transport of solutes because they flow with, i.e., are convected with, the water that flows through the membrane (first term), and transport because of a solute concentration gradient (second term). Here, x is a coordinate axis directed across the membrane. The diffusional term depends on a membrane-based diffusion coefficient, $D_{\text{m},i}$, which includes the effects of porosity and tortuosity. The friction factor $K_{\text{f},i}$ depends on the extent of solute-membrane friction, and thus relates to membrane pore size and the (hydrated) size of the solutes. In the absence of solute-membrane friction, the solutes move freely with the water in the pores, while they are still subject to diffusional forces (and electrical forces when charged), and in that case $K_{\text{f},i}\!=\!1$. However, in an RO process, it is much more likely that $K_{\text{f},i}\ll 1$.
@@
$\Phi_{i}=\frac{c_{\text{m},i}}{c_{\infty,i}}$ (14)
where $\Phi$ is a partition coefficient; index ‘m’ refers to a position just in the membrane and $\infty$’ to a position just outside the membrane. When the solutes are charged, i.e., they are ions, an additional Donnan effect arises, which in this paper we treat separately, i.e., do not include in $\Phi_{i}$, see section 5.2. The partition coefficient, $\Phi_{i}$, can be due to the chemical affinity of a solute with a certain phase (medium) relative to its affinity with another phase, but also other effects of solute-membrane interaction can be absorbed in $\Phi_{i}$ [23, 24, 17]. In RO with neutral solutes, we always have at least $\Phi_{i}\!<\!1$ or $K_{\text{f},i}\!<\!1$, and likely both factors are $<\!1$.
A further element in RO modeling is the permeate equation (Eq. (8) in [5]; Eq. (24) in [25])
$c_{\text{p},i}=\frac{J_{i}}{J\mathrm{{}_{w}}}$ (15)
which predicts the concentration of solutes on the permeate side. It a cocurrent flow geometry, it is only valid when at each point in the module the concentration on the permeate side is the same. Thus it is not generally valid when the permeate water flows along the membrane to a single exit point. But for a module in the limit of a very small water recovery, or when on the upstream side we have a mixed solution (stirred tank), it is valid. In any module it describes the concentration in the exit of the permeate flow, when we use average values for $J_{i}$ and $J\mathrm{{}_{w}}$ (each separately averaged over the complete module). It is also valid in cross-current flow for certain conditions [14].
@@
${J_{i}}=\left(1-\sigma_{i}\right)\,J\mathrm{{}_{w}}\,\frac{c_{\text{f},i}\cdot e^{\text{Pe}_{i}}-c_{\text{p},i}}{e^{\text{Pe}_{i}}-1}$ (16)
where $\text{Pe}_{i}$ is the membrane Péclet number, $\text{Pe}_{i}=J\mathrm{{}_{w}}/k_{\text{m},i}$, where $k_{\text{m},i}$ is the membrane mass transfer coefficient, given by $k_{\text{m},i}=D_{\text{m},i}{\big{/}}L\mathrm{{}_{m}}$, with Lm membrane thickness. In Eq. (16), $\sigma_{i}$ is the sieving coefficient (also called reflection coefficient) which relates to $K_{\text{f},i}$ and $\Phi_{i}$ according to
@@
Eq. (16) is the general result for rejection in an RO membrane with neutral solutes [25, 29, 27, 28, 26, 5]. Calculation results based on this equation are presented in Fig. 3. Note that when concentration polarization (CP) is also considered, Eq. (16) is still valid, but we must replace $c_{\text{f},i}$ by the concentration right at the membrane interface, often written as $c_{\text{int},i}$.
@@
which is equivalent to Eq. (4) in ref. [21], Eq. (3) in ref. [28], Eq. (38) in ref. [30], and Eq. (129) in ref. [31], when we implement that $\omega=\left(1-\sigma_{i}\right)\,k_{\text{m},i}$, $P=P\mathrm{{}_{s}}=K_{\text{f}}\Phi_{i}k_{\text{m},i}$, and thus $k_{\text{m},i}=P\mathrm{{}_{s}}/\left(1-\sigma_{i}\right)$. A similar formulation is Eq. (9) in ref. [32]. Note that while Eq. (16) is generally valid, at each position in an RO module, Eq. (18) makes use of the permeate equation, Eq. (15). All equations in this section are derived for neutral solutes.
Based on Eq. (18), rejection in the limit of low transmembrane water flux $J\mathrm{{}_{w}}$, is given by
$R_{i}=\frac{\sigma_{i}}{1-\sigma_{i}}\;\text{Pe}_{i}=\frac{\sigma_{i}}{\omega}\;J\mathrm{{}_{w}}$ (19)
which shows that for $J\mathrm{{}_{w}}\rightarrow 0$, solute rejection goes to zero, and it also shows that to have a non-zero rejection we need $\sigma_{i}>0$, and thus either $K_{\mathrm{f},i}$ must be $<1$, or $\Phi_{i}$ must be $<1$.
In the other limit, of a high permeate water flux, $J\mathrm{{}_{w}}$, Eq. (16) simplifies to
$J_{i}=K_{\text{f},i}\Phi_{i}\,c_{\text{f},i}\,J\mathrm{{}_{w}}=\left(1-\sigma_{i}\right)\,c_{\text{f},i}\,J\mathrm{{}_{w}}$ (20)